# Thin films spreading under the action of surface acoustic waves

### v1

@@ Abstract
Experimentally and theoretically, we consider the problem of a spreading drop of millimetric thickness driven by a surface acoustic wave (SAW). The experimental results show that, after initial transients, the drop spreads with approximately constant speed while leaving a thin layer of fluid behind. To better understand the system, we develop a theoretical model that couples acoustic driving with the fluid mechanics of a spreading drop. An important feature of the problem is that the strength of the acoustic driving depends on the fluid thickness, with strong attenuation of the forcing in regions of the domain covered by fluid. The presented model provides results consistent with the experiments, particularly regarding drop height and spreading speed. The constant spreading speed observed in the experiments is also captured by an approximate traveling-wave solution, which provides predictions for the dependence of drop thickness and spreading speed on the acoustic amplitude.
1 Introduction
@@
Dynamic wetting of a solid substrate by a film of liquid is a common occurrence. In natural systems it appears in surfactant films that wet the lungs to promote breathing and the eyes to keep them moist (Veldhuizen & Haagsman, 2000; Holly & Lemp, 1971), or in the spreading of water drops on solid surfaces (such as raindrops on windows) (Dussan V. & Davis, 1974; Dussan V. & Chow, 1983). In technological systems it is harnessed for actuating microfluidic platforms (Atencia & Beebe, 2005; Stone et al., 2004; Whitesides, 2006), for cooling electronic circuits (Amon et al., 2001; Bar-Cohen et al., 2006), for desalination (Fletcher et al., 1974) and for a variety of manufacturing applications (Fendler, 1996; Nagayama, 1996; Wang et al., 2004). Many different mechanisms govern dynamic wetting, where capillary forces at the free surface of a liquid film or a meniscus contribute appreciably to the dynamics of the film. In addition to capillary forces, dynamic wetting may be powered by gravity, thermal or solutal Marangoni effects, or electrokinetics, among others, see Oron et al. (1997); Craster & Matar (2009) for comprehensive reviews. In the present work we focus on the contributions of mechanical vibrations in the solid substrate, particularly MHz-frequency Rayleigh surface acoustic waves (SAWs), to dynamic wetting (Morozov & Manor, 2018). Here, we concentrate on the contribution of Eckart streaming – mass transport that results from the attenuation of sound or ultrasound waves in the bulk of the liquid film. While extensively observed and described in the scientific literature (Shiokawa et al., 1989; Brunet et al., 2007, 2009; Yeo & Friend, 2014), dynamic wetting of a solid substrate by a liquid film, induced by Eckart streaming, was not studied in detail to date. In this work we use thin film (long wave) theory alongside physical experiments to elucidate the phenomenon.
@@
There are three types of acoustic streaming mechanisms by which acoustic waves or waves at the free surface of a liquid or a solid boundary (surface acoustic waves) generate streaming – steady flow along the path of the wave. Firstly, a wave – acoustic wave or surface wave – in the solid or in liquid, invokes periodic viscous flow within the viscous penetration length $\delta\equiv\sqrt{2\mu/\rho\omega}$ away from the interface, where $\mu,~{}\rho$ and $\omega$ are the viscosity and density of the liquid and the wave angular frequency, respectively. Examples in which this problem was considered include the work by Rayleigh on a standing acoustic wave grazing a solid bed; by Schlichting (1932) on standing vibration of waves of infinite and finite (respectively) wavelength in a solid along its surface; by Longuet-Higgins (1953) on shallow ocean waves; and by Manor et al. (2012) on propagating Rayleigh (surface acoustic) waves in a solid substrate. Convective contributions due to surface waves invoke drift of liquid mass, which does not attenuate away from the solid surface and is known variously as the Stokes drift (Stokes, 1847), the Rayleigh law of streaming (Lighthill, 1978), and Schlichting streaming (Yeo & Friend, 2014); although the physics and mathematical modeling of the cases studied by Stokes, Rayleigh, and Schlichting are similar, Rayleigh is the only one among them that explicitly considered acoustic waves.
Secondly, in the presence of an acoustic wave in the liquid, whether due to an actuator or an acoustic wave in the solid, one should expect acoustic radiation pressure (Campbell, 1970; Shiokawa et al., 1989). Convective effects – Reynolds stresses – within the wave increase the local steady pressure in the bulk of the liquid; upon reflection of the wave at an interface between phases of mismatched acoustic impedance (density multiplied by acoustic phase velocity), the wave imposes convective stress, known as the acoustic radiation pressure (King, 1934; Borgnis, 2010; Hasegawa et al., 2000), which is capable of deforming and displacing the interface (Alzuaga et al., 2005; Biwersi et al., 2000; Rajendran et al., 2022).
Finally, a third (and the most well-known) mechanism that may appear due to the attenuation of sound or ultrasound waves in fluid is Eckart streaming (Eckart, 1948; Lighthill, 1978; Nyborg, 1952), where wave intensity is converted to fluid momentum, characterized by an intense vortical flow field that appears in the bulk fluid. This is the main acoustic mechanism for the actuation of fluid in micro-channels (Yeo & Friend, 2014; Wixforth et al., 2004; Wixforth, 2003) and in drop microfluidics (Guttenberg et al., 2004; Brunet et al., 2007, 2009, 2010).
These different acoustofluidic mechanisms govern mass transport at different length scales (Rezk et al., 2014). In the presence of sound or ultrasound waves in the fluid, the surface effect, i.e., Stokes drift, governs the transport of mass at separations from the solid that are smaller than the acoustic wavelength in the fluid. Further away, mass transport in the fluid is governed by Eckart streaming. Acoustic radiation pressure effects may be relevant in the former case, and also in the latter case when Eckart streaming is weak. Moreover, SAW diffraction adds an additional spatial dimension to flow actuation by SAW. While the SAW wavelength usually determines the lower spatial resolution of the resulting acoustic flow, SAW diffraction allows for breaking the SAW wavelength limit as a measure for spatial flow resolution (Fakhfouri et al., 2018).
@@
There are abundant studies on the dynamic wetting of thin liquid films that are governed by Stokes drift and acoustic radiation pressure. Rayleigh surface acoustic waves (SAWs) were recently shown to power the continuous spreading of sub-micrometer to hundreds of micrometer thick films of fully wetting silicon oils (Rezk et al., 2012, 2014) and partially wetting water/surfactant mixtures (Altshuler & Manor, 2015, 2016; Mhatre et al., 2016) along or opposite the path of the propagating SAW; water mixtures evaporate leaving behind SAW-induced deposit morphology. However, we are not aware of studies that consider in detail the acoustic/capillary balance of dynamic wetting in liquid films that are sufficiently thick that mass transport therein, and hence film dynamics, are actuated by Eckart streaming.
@@
In this paper, we use physical experiments, theoretical modeling and simulations to elucidate the physics of dynamic wetting powered by Eckart streaming under the action of MHz-frequency Rayleigh-type SAWs. We describe our experimental findings in Sec. 2 and then proceed with describing the theoretical model in Sec. 3. Theoretical results are presented in Sec. 4; we first consider computational results in Secs. 4.1 and 4.2 and then proceed with an exploration of translationally-invariant solutions in Sec. 4.3. Section 5 is devoted to the summary and conclusions.
2 Experiment
[…]
@@ 2 Experiment
Figure 1: (A) Upper schematic view of the experimental setup: Surface acoustic wave (SAW) propagates from the interdigitated transducer (IDT) until it reaches the acoustic absorber (comprised of glycerol-soaked paper placed on the actuator under and to the right of the needle). The needle is of known diameter 510 $\mu$m, placed for identifying spatial resolution in the images. (a-e) Successive snapshots taken from an experiment monitoring the flow of a silicone oil film. During the experiment, a drop of silicone oil is placed on the horizontal surface (a), deforms due to the application of SAW (b), and moves in the direction of SAW propagation (c-e).
[…]
@@
Figure 2: Two sets of images from experiments monitoring silicone oil powered by SAWs of three different amplitudes (1.2 nm-1.4 nm), where (A-C) silicone oil starts from the same location, moves in the direction of the SAW, and reaches different distances (a-c) after 4.09 seconds. The corresponding side-view videos at different acoustic power levels are used to analyze the speed and profile of silicone oil during the movement. (D, d) show top view snapshots (corresponding to (C,c)) of the oil. A thin film of oil, hardly visible in the side view, can be seen more clearly behind the main body of silicone oil in the top views.
@@
Figure 4: (a) Time evolution of the front position for several SAW amplitudes for $A=0.52$ nm (black), $0.65$ nm (gray), $0.78$ nm (purple), $0.91$ nm (cyan), $1.04$ nm (green), $1.17$ nm (magenta), $1.30$ nm (red), $1.43$ nm (brown), $1.56$ nm (orange), $1.69$ nm (blue). The arrow points the direction of increasing $A$. (b) Time variations of the maximum drop height for SAW amplitudes $A=0$ nm (green, no SAW applied), $0.5$ nm (blue), $0.7$ nm (orange), $1.3$ nm (yellow) and $1.6$ nm (purple). The symbols correspond to the measurements and the lines are simple fits to guide the eye. We expect errors of ±$\pm 50~{}\mu$m in the values of the film height due to the limited resolution of the side view image and tracker software.
@@
In our experiments, we vary the SAW displacement amplitude at the surface of the actuator by varying the applied voltage level input to the SAW device (there is a linear relationship between the two; see Appendix A for the corresponding calibration graph). To confirm that the observed dynamics are indeed induced by SAW, we also carried out control experiments, shown in Fig. 3, in which SAW is absent and all other parameters are kept the same. The oil puddle in this experiment, which was repeated three times, slowly spreads concentrically, as expected in the absence of a directional forcing mechanism.
@@
We then employ the side view video, recorded by a goniometer camera, to capture the position of the advancing three-phase contact line of the SAW-actuated oil film using the open-source software Tracker (Brown & Cox, 2009) (see Fig. 4(a) for typical data). We also measure the maximum height of the moving oil film using the public domain image processing software ImageJ (typical results in Fig. 4(b)). Figure 4(a) shows that, as expected, the larger the SAW amplitude $A$, the faster the oil film moves due to the stronger induced leakage wave (Shiokawa et al., 1989). Another notable aspect of SAW-induced dynamics is the contact line speed for early times ($<1$ s): initially, it is rather small, and later it increases and then remains constant. This change of contact line speed appears to be a consequence of the initial change of shape of the drop (see also Fig. 1); while this is taking place, contact line speed is small. After the initial deformation, the drop shape remains unchanged, translating approximately uniformly (this will be seen in Fig. 9 later in the text). Figure 4(b) shows consistent results for the maximum drop height, with an increase of the height for early times, and constant values for longer times. Note that this uniform translation and constant film height differ significantly from the observed behavior for spreading under a body force such as gravity (for gravity-driven spreading such as in Fig. 3 the contact line speed decreases, and the drop becomes thinner as it spreads). These experimental observations will be explained and discussed within the framework of our theoretical model, which is described next.
3 Model
@@
To model the experiments described in Sec. 2, we consider a fluid traversed by a sound wave. The fluid velocity, $\mathbf{V}$, can be written as
@@
where $\mathbf{v}$ is the “bulk” velocity of the fluid, associated with a characteristic time of the flow, $t_{c}$, which is on the order of seconds (so $|\mathbf{v}|\sim\ell/t_{c}$, if $\ell$ is a typical lengthscale of the flow), and $\mathbf{v}_{s}$ is associated with the much faster timescale of the sound wave, $T=f^{-1}\ll t_{c}$ (typically $T\sim 10^{-6}$ s). The idea behind Eq. (1) is that $\mathbf{v}_{s}$ corresponds to an oscillatory flow generated by the propagating sound wave, while $\mathbf{v}$ is the velocity due to external forces (or to the wave propagation itself, as we will show below). Therefore, the fast-time averages, $\langle\cdot\rangle=(1/T)\int_{0}^{T}\cdot\,dt$, of $\mathbf{v}$ and $\mathbf{v}_{s}$, are given by
@@
We write now the Navier–Stokes (NS) equations for the fluid in the compact form
@@
and $\rho_{0}=\langle\rho\rangle$ is the average (constant) density of the fluid (note that $\langle\partial(\rho_{0}\mathbf{v}_{s})/\partial t\rangle=0$). Equation (4) shows that the flow field $\mathbf{v}$, which develops on the long time scale $t_{c}$, satisfies NS equations similar to those for $\mathbf{V}$, but with an additional volume force, $\mathbf{F}_{s}$. This flow field, $\mathbf{v}$, is known as acoustic streaming flow, resulting from the high-frequency sound wave traversing the fluid (see, e.g. Shiokawa et al. (1989)).
When the sound wave is a leaky wave generated at a solid substrate that propagates in the $(x,z)$–plane (see Fig. 5), the fluid particle displacement, $\mathbf{u}$, can be written as (Shiokawa et al., 1990)
@@
where $A$ is the wave amplitude, $k_{L}=k_{r}+\imath\,k_{i}$ is the complex wavenumber of the leaky wave and $x^{\prime}=x-x_{\rm ref}(t)$, where $x_{\rm ref}(t)$ is a reference $x$–coordinate determining the onset of attenuation, whose origin will be discussed in Sec. 4. Also, we define
$\alpha_{1}=\sqrt{(c_{s}/c_{\rm oil})^{2}-1},$ (9)
where $c_{s}$ and $c_{\rm oil}$ are the SAW phase velocities in the substrate and silicon oil, respectively. By choosing (Roditi, 2023) $c_{s}=3880$ m/s and (PrecisionFlow, 2023) $c_{\rm oil}=1350$ m/s, we obtain that $\alpha_{1}=2.69$. Thus, the Rayleigh angle, $\theta$, is given by $\theta=\arctan\alpha_{1}=69.65^{\circ}$. The SAW attenuation factor is given by (Arzt et al. (1967))
$k_{i}=\frac{1}{l_{i}}=\frac{\rho_{\rm oil}}{\rho_{s}}\frac{c_{\rm oil}}{c_{s}^{2}}f,$ (10)
[…]
Figure 5: Schematic of a liquid drop (blue curve) forced by a SAW traveling from left to right, with rear and front contact lines at $x_{r}(t)$ and $x_{f}(t)$, respectively. The thin red line represents the amplitude $A$ of the leaky SAW at the substrate and shows how it is attenuated by liquid (the black dashed line is the envelope of the decaying amplitude in the liquid). The Rayleigh angle is denoted by $\theta$.
@@
which, on substitution into Eq. (5), leads to explicit expressions for the force components,
${\rm F}_{s,x}$ $=$ $\left(1+\alpha_{1}^{2}\right)k_{i}\rho_{0}A^{2}\omega^{2}e^{-2k_{i}\left(x^{\prime}+\alpha_{1}z\right)},$ (12)
@@
$\phi_{s}(x,z)=\phi_{0}\,e^{-2k_{i}\left(x^{\prime}+\alpha_{1}z\right)},\quad\mbox{with}\quad\phi_{0}=\frac{1+\alpha_{1}^{2}}{2}\rho_{0}A^{2}\omega^{2}.$ (15)
Therefore, the NS equations for the velocity field $\mathbf{v}$, under the influence of both SAW and gravitational forces, can be written as
$\rho_{0}\frac{\partial\mathbf{v}}{\partial t}=-\nabla P+\mu\nabla^{2}\mathbf{v},$ (16)

### v2

# Dynamics of a millimeter size drop exposed to MHz-frequency surface acoustic wave in the underlying solid substrate

@@ Abstract
We present an experimental, theoretical, and computational study of the spreading of a millimetric silicon oil drop under the excitation of a surface acoustic wave (SAW), traveling in the underlying solid substrate. Our fully time-dependent theoretical model is formulated within the long wave approach, and accounts carefully for the interaction between fluid dynamics and acoustic driving. While an approach similar to ours was used in previous works to analyze the dynamics of micron-scale films of oil and water under SAW excitation, the acoustic forcing was associated with boundary layer flow, i.e., Schlichting and Rayleigh streaming, and with acoustic radiation pressure. For the case considered here (millimetric thickness drops), the acoustic forcing is related to the Eckart streaming: the forcing mechanism is a result of variations in the Reynolds stress in the liquid, which is associated with variations in the intensity of ultrasonic wave leakage off the attenuating SAW under the drop. Schlichting and Rayleigh streaming contributions are usually negligible in this case. We find, in both experiments and simulations, that following an initial transient, where the oil drop deforms to accommodate the acoustic stress therein, the drop accelerates to reach translation at approximately constant speed for long times, while leaving a thin wetting layer behind. Our model suggests that the drop’s steady speed is due to the quasi-steady shape of the drop’s main body at long times, which is itself a consequence of the rate of SAW attenuation, which depends on the thickness of the liquid body above. The SAW attenuation is weak under air and the micron-thick liquid trail left behind the drop; it becomes significant once the SAW reaches the main body of the drop. Hence, the forcing on the drop is approximately the same, irrespective of its distance to the source of the SAW. The translating drop’s speed is found to be dependent on the drop size, as well as on the SAW intensity. The drop’s steady shape and speed are further informed by analytical consideration of a simplified traveling wave type model, that clarifies contributions from different physical effects. The theoretical results show qualitative and sometimes even quantitative agreement with the experiments, particularly regarding drop height and traveling speed.
I Introduction
@@
Dynamic wetting of a solid substrate by a liquid film is a common occurrence. In natural systems, it appears in surfactant films that wet the lungs to promote breathing and the eyes to keep them moist [2, 1], or in the spreading of water drops on solid surfaces (such as raindrops on windows) [4, 3]. In technological systems it is harnessed for actuating microfluidic platforms [5, 6, 7], for cooling electronic circuits [8, 9], for desalination [10] and for a variety of manufacturing applications [11, 12, 13]. Many different mechanisms govern dynamic wetting, where capillary forces at the free surface of a liquid film or a meniscus contribute appreciably to the dynamics of the film. In addition to capillary forces, dynamic wetting may be powered by gravity, thermal or solutal Marangoni effects, or electrokinetics, among others, see Oron et al. [14] or Craster and Matar [15] for comprehensive reviews. In the present work, we focus on the contributions of mechanical vibrations in the solid substrate, particularly MHz-frequency Rayleigh surface acoustic waves (SAWs), to dynamic wetting [16]. We study particularly the effects of Eckart streaming – mass transport in the bulk fluid that results from variations in acoustic or ultrasonic intensity. While extensively observed and described in the scientific literature [17, 18, 19], we are not aware of a model describing the dynamics of such a system that further accounts for the capillary and gravitational stresses that shape the liquid/air interface for the case when drops are sufficiently thick (millimeter scale). Here, we use thin film (long wave) theory alongside physical experiments to develop a better understanding of the dynamics of this problem.
@@
There are several different mechanisms by which acoustic waves or other mechanical waves, at a solid boundary or at the free surface of a liquid, generate streaming – steady flow along the wave’s path. A mechanical wave at a boundary invokes periodic viscous flow within the viscous penetration length $\delta\equiv\sqrt{2\mu/(\rho\omega)}$ away from the interface, where $\mu,~{}\rho$ and $\omega$ are the viscosity and density of the liquid, and the wave angular frequency, respectively. Examples in which this problem was considered include the work by Rayleigh on a standing acoustic wave grazing a solid bed [20]; by Schlichting [21] on standing vibration of waves of infinite and finite (respectively) wavelength in a solid along its surface; by Longuet-Higgins [22] on shallow ocean waves; and by Manor et al. [23] on propagating Rayleigh (surface acoustic) waves in a solid substrate. Convective contributions due to surface waves invoke a drift of liquid mass, which does not attenuate away from the solid surface. It is known as Schlichting streaming [19], and its component far from the boundary was further coined the “Rayleigh law of streaming” by Lighthill [24] in recognition of Rayleigh’s work on the drift resulting from the presence of standing acoustic waves in fluid near a solid boundary.
Another mechanism, the Eckart streaming – the focus of the current work – appears regardless of the presence of a boundary. Eckart streaming results from variations of the intensity of sound or ultrasound waves in the fluid. In the classic work by Eckart and in many subsequent studies, the Eckart streaming is ascribed to the viscous attenuation of traveling sound waves in the fluid [25, 24, 26]. This attenuation results in spatial variations of the fluid velocity field, and therefore in the convective Reynolds stresses. The consequence is the occurrence of a secondary flow whose steady component at long times is the acoustic streaming. Eckart streaming is characterized by an intense vortical flow field that appears in the bulk fluid. This is the main acoustic mechanism for the actuation of fluid in micro-channels [19, 28, 27] and in drop microfluidics [30, 17, 18, 29]. It usually dominates contributions to the flow in the bulk fluid arising from the boundary layer type acoustic streaming mechanisms (such as Shlichting or Rayleigh streaming described above).
Approximately 40 years after Eckart’s work, Shiokawa et al. [31] considered MHz-frequency SAWs that travel in a solid and leak same-frequency ultrasonic waves into the fluid. In this scenario, they showed that the main contribution to spatial variations in wave intensity in the fluid and, hence, to variations in the Reynolds stress arises from the fast attenuation of surface acoustic waves (SAWs) in the neighboring solid substrate. Moreover, the SAW attenuation is faster than that of the viscous attenuation of the ultrasound leakage waves as they travel in the fluid. Thus, to leading order, variation in the intensity of ultrasound leakage in the fluid results from SAW attenuation in the solid. For example, 20 MHz frequency SAWs traveling in a solid substrate in contact with a thick layer of water at ambient conditions possess an attenuation length of approximately 2 millimeters [32]. However, the same frequency ultrasonic waves in bulk water possess an attenuation length of approximately 100 millimeters [33]. Hence, in contrast to the assumptions in Eckart’s original work on acoustic streaming, the Reynolds stress variation in a fluid in contact with a leaky SAW is to leading order correlated with the attenuation of the SAW in the solid, rather than with the viscous attenuation of the ultrasound wave leakage in the fluid.
@@
In addition to the flow mechanisms discussed above, the presence of an acoustic wave in a liquid, whether due to an actuator or an acoustic wave in the solid, leads to acoustic radiation pressure [34, 35, 31, 36, 37]. Convective effects (Reynolds stresses) within the wave increase the local steady pressure in the bulk of the liquid; upon reflection of the wave at an interface between phases of mismatched acoustic impedance (density multiplied by acoustic phase velocity), the wave imposes convective stress – acoustic radiation pressure [40, 38, 39] – capable of deforming and displacing the interface [41, 42, 43]. There are abundant studies on the dynamic wetting of thin liquid films that are governed by Schlichting and Rayleigh streaming, alongside acoustic radiation pressure. For example, SAWs were recently shown to power the continuous spreading of sub-micrometer to tens of micrometers thick films of fully wetting silicon oils [44, 45] and partially wetting water/surfactant mixtures [46, 47, 48] along or opposite the path of the propagating SAW. However, we are unaware of studies that consider the acoustic/capillary balance of dynamic wetting in liquid films and drops that are sufficiently thick that mass transport therein, and hence film dynamics, are actuated by Eckart streaming. Consideration of such a regime is the subject of the present paper.
@@
In this paper, we use theory and simulations, which we compare to our physical experiments, to elucidate the physics of dynamic wetting powered by Eckart streaming under the action of MHz-frequency Rayleigh-type SAWs. We describe our experiments in Sec. II and then derive the theoretical model in Sec. III. Our theoretical results are presented in Sec. IV: we discuss computational results in Secs. IV.1 and IV.2, before analyzing a simplified traveling–wave type model in Sec. IV.3. Section V is devoted to the summary and conclusions.
II Experiment
Figure 1 shows the experimental setup. We generate a propagating 20 MHz frequency surface acoustic wave (SAW) by applying a same-frequency sinusoidal voltage signal to a piezoelectric actuator – a SAW device. The actuator comprises a 5 nm titanium/1 $\mu$m aluminum interdigitated transducer (IDT; from which the SAW emanates) fabricated atop lithium niobate (LiNbO3, Roditi International, UK) by standard lift-off photolithography. The substrate used for the SAW device is 11 mm $\times$ 24 mm in size, 0.5 mm thick, 128° Y-cut, X-propagating, single-crystal piezoelectric lithium niobate, where X and Y are crystal axes [49]. The actuator is integrated into the external electrical signal using pogo pins (BC201403AD, Interconnect Devices, Inc.) assembled in a 3D-printed elastomeric stage, which holds the actuator and is connected to a signal generator (R&S SMB100A microwave signal generator) and amplifier (model A10160, Tabor Electronics Ltd.). We place a paper cylinder soaked with glycerol at the far end of the actuator from the IDT. The glycerol-soaked paper absorbs the SAW and prevents reflections. We measure the SAW normal surface vibration at the actuator surface and verify that it is a propagating wave using a scanning laser Doppler vibrometer (MSA-500, Polytech).
@@ II Experiment
Figure 1: (A) Upper schematic view of the experimental setup: Surface acoustic wave (SAW) propagates from the interdigitated transducer (IDT) until it reaches the acoustic absorber (comprised of glycerol-soaked paper placed on the actuator under and to the right of the needle). The needle is of known diameter (510 $\mu$m), placed for identifying spatial resolution in the images. (a-e) Successive snapshots are taken from an experiment monitoring the flow of a silicone oil film. During the experiment, a drop of silicone oil is placed on the horizontal surface (a), it deforms due to the application of SAW (b), and moves in the direction of SAW propagation (c-e); time is shown in seconds, and the vertical lines serve as a reference.
We place an 8 mm3 ($8~{}\mu$l) puddle of silicone oil (50 cSt, 378356, Sigma-Aldrich) atop the actuator using a pipette, approximately 3.5 mm away from the IDT, and introduce electrical signal at different voltage levels to induce motion of the oil film, which dynamically wets the solid substrate along the path of the SAW. Figure 2(a-c) shows a time-lapse (side view) of the silicone oil dynamics, captured using the framework and camera of a goniometer (Data Physics; OCA 15Pro). Figure 2(d) shows a top view obtained using a camera (EOS R5, Canon) with a macro lens (RF 100 mm F2.8L MACRO IS USM, Canon). Prior to the experiment we clean the SAW device using four different solvents: acetone (AR-b, 99.8%, 67-641, Bio-Lab Ltd.), 2-propanol (AR-b, 99.8%, 67-63-0, Bio-Lab Ltd.), ethanol (CP-p, 96%, 64-17-5, Bio-Lab Ltd.), and water (HPLC plus, 7732-18-5, Sigma-Aldrich). Finally, we dry the actuator using compressed air.
@@
Figure 2: Two sets of images from experiments monitoring silicone oil powered by SAWs of three different amplitudes $A$ (1.2 nm–1.4 nm), where (A-C) silicone oil starts from the same location, moves in the direction of the SAW, and reaches different distances (a-c) after 4.09 seconds. The corresponding side-view videos at different acoustic power levels are used to analyze the speed and profile of silicone oil during the movement. (D, d) show top view snapshots (corresponding to (C,c)) of the oil. A thin film of oil, hardly visible in the side view, can be seen more clearly behind the main body of silicone oil in the top views.
@@
Figure 4: (a) Time evolution of the front position for several values of $A=0.52$ nm (black), $0.65$ nm (gray), $0.78$ nm (purple), $0.91$ nm (cyan), $1.04$ nm (green), $1.17$ nm (magenta), $1.30$ nm (red), $1.43$ nm (brown), $1.56$ nm (orange), $1.69$ nm (blue). (b) Time variations of the maximum drop height for $A=0$ nm (gray, no SAW applied), $0.52$ nm (black), $0.78$ nm (purple), $1.30$ nm (red) and $1.69$ nm (blue). The symbols correspond to the measurements and the lines are simple fits to guide the eye. We expect errors of ±$\pm 50~{}\mu$m in the values of the film height due to the limited resolution of the side view image and tracker software.
@@
In our experiments, we vary the SAW displacement amplitude $A$ (normal amplitude at the solid surface before any attenuation) at the surface of the actuator by varying the applied voltage level input to the SAW device (there is a linear relationship between the two; see Appendix A for the corresponding calibration graph). To confirm that the observed dynamics are indeed induced by SAW, we also carried out control experiments, shown in Fig. 3, in which SAW is absent and all other parameters are kept the same. The oil puddle in this experiment, which was repeated three times, slowly spreads concentrically, as expected in the absence of a directional forcing mechanism.
@@
We employ the side view video, recorded by a goniometer camera, to capture the position of the advancing three-phase contact line of the SAW-actuated oil film using the open-source software Tracker [50] (see Fig. 4(a) for typical data). We also measure the maximum height of the moving oil film using the public domain image processing software ImageJ (typical results in Fig. 4(b)). Figure 4(a) shows that, as expected, the larger the SAW amplitude $A$, the faster the oil film moves due to the stronger induced leakage wave [31]. Another notable aspect of SAW-induced dynamics is the contact line speed for early times ($<1$ s): initially, it is rather small, but later increases and then remains constant. This change of contact line speed appears to be a consequence of the initial change of shape of the drop (see also Fig. 1); while this is taking place, contact line speed is small. After the initial deformation, the drop shape remains unchanged, translating approximately uniformly (this will be seen in Fig. 9 later in the text). Figure 4(b) shows consistent results for the maximum drop height, with an increase of the height for early times, and constant values for longer times. Note that this uniform translation and constant film height differ significantly from the observed behavior for spreading under a body force such as gravity (for gravity-driven spreading such as in Fig. 3 the contact line speed decreases, and the drop becomes thinner as it spreads). These experimental observations will be explained and discussed within the framework of our theoretical model, which is described next.
III Model
@@
To model the experiments described in Sec. II, we consider a fluid traversed by a sound wave. The fluid velocity, $\mathbf{V}$, can be written as
@@
where $\mathbf{v}$ is the “bulk” velocity of the fluid, associated with a characteristic time of the flow, $t_{c}$, which is on the order of seconds (so $|\mathbf{v}|\sim\ell/t_{c}$, if $\ell$ is a typical lengthscale of the flow); and $\mathbf{v}_{s}$ is associated with the much faster timescale of the sound wave, $T=f^{-1}\ll t_{c}$ (typically $T\sim 10^{-6}$ s). The idea behind Eq. (1) is that $\mathbf{v}_{s}$ corresponds to an oscillatory flow generated by the propagating sound wave, while $\mathbf{v}$ is the velocity due to external forces (or to the wave propagation itself, as we will show below). Therefore, the fast-time averages, $\langle\cdot\rangle=(1/T)\int_{0}^{T}\cdot\,dt$, of $\mathbf{v}$ and $\mathbf{v}_{s}$, are given by
@@
We write the Navier–Stokes (NS) equations for the fluid in the compact form
@@
and $\rho_{0}=\langle\rho\rangle$ is the average (constant) density of the fluid (note that $\langle\partial(\rho_{0}\mathbf{v}_{s})/\partial t\rangle=0$). Equation (4) shows that the flow field $\mathbf{v}$, which develops on the long time scale $t_{c}$, satisfies NS equations similar to those for $\mathbf{V}$, but with an additional volume force, $\mathbf{F}_{s}$. This flow field, $\mathbf{v}$, is known as acoustic streaming flow, resulting from the high-frequency sound wave traversing the fluid (see, e.g. [31]). As we pointed out in the Introduction, this acoustic streaming flow corresponds to the Eckart streaming due to the Reynolds stress variation in the fluid caused by SAW attenuation in the solid.
When the sound wave is a leaky wave generated at a solid substrate that propagates in the $(x,z)$–plane (see Fig. 5), the fluid particle displacement, $\mathbf{u}$, can be written as [51] (we discuss later in this section the assumptions behind this formulation)
@@
where $A$ is the wave amplitude, $k_{L}=k_{r}+\imath\,k_{i}$ is the complex wavenumber of the leaky wave and $x^{\prime}=x-x_{\rm ref}(t)$, where $x_{\rm ref}(t)$ is a reference $x$–coordinate determining the onset of attenuation, whose origin will be discussed in Sec. IV. Also, we define
$\alpha_{1}=\sqrt{(c_{\rm s}/c_{\rm oil})^{2}-1},$ (9)
where $c_{\rm s}$ and $c_{\rm oil}$ are the SAW phase velocities in the substrate and silicon oil, respectively. By choosing [52] $c_{\rm s}=3880$ m/s and [53] $c_{\rm oil}=1350$ m/s, we obtain $\alpha_{1}=2.69$. Thus, the Rayleigh angle $\theta$ is given by $\theta=\arctan\alpha_{1}=69.65^{\circ}$. The SAW attenuation factor is given by [32]
$k_{i}=\frac{1}{l_{i}}=\frac{\rho_{\rm oil}}{\rho_{\rm s}}\frac{c_{\rm oil}}{c_{\rm s}^{2}}f,$ (10)
where $l_{i}$ is the attenuation length (of the SAW in the solid), and $\rho_{\rm s}$ is the solid density. For the experimental values: $f=20$ MHz, $\rho_{\rm s}=4.65$ g/cm3 and $\rho_{\text{oil}}=0.96$ g/cm3, we obtain $l_{i}=2.7$ mm. Note that the SAW wavelengths in the solid substrate and the oil are $\lambda_{\rm s}=c_{\rm s}/f=194$ $\mu$m and $\lambda_{\rm oil}=c_{\rm oil}/f=67.5$ $\mu$m, respectively.
Figure 5: Schematic of a liquid drop (blue curve) forced by a SAW traveling from left to right, with rear and front contact lines at $x_{\rm r}(t)$ and $x_{\rm f}(t)$, respectively. The thin red line represents the amplitude $A$ of the leaky SAW at the substrate and shows how it is attenuated by liquid (the black dashed line is the envelope of the decaying amplitude in the liquid). The Rayleigh angle is denoted by $\theta$.
@@
which, on substitution into Eq. (5), leads to explicit expressions for the force components (note that we use $\rho_{0}=\rho_{\rm{oil}}$ from now on to avoid confusion),
${\rm F}_{s,x}$ $=$ $\left(1+\alpha_{1}^{2}\right)k_{i}\rho_{\rm{oil}}A^{2}\omega^{2}e^{-2k_{i}\left(x^{\prime}+\alpha_{1}z\right)},$ (12)
@@
$\phi_{s}(x,z)=\phi_{0}\,e^{-2k_{i}\left(x^{\prime}+\alpha_{1}z\right)},\quad\mbox{with}\quad\phi_{0}=\frac{1+\alpha_{1}^{2}}{2}\rho_{\rm{oil}}A^{2}\omega^{2}.$ (15)
Therefore, the NS equations (4) for the velocity field $\mathbf{v}$, under the influence of both SAW and gravitational forces, can be written as
$\rho_{\rm{oil}}\frac{\partial\mathbf{v}}{\partial t}=-\nabla P+\mu\nabla^{2}\mathbf{v},$ (16)